Trying to prove this is an identity.
cos^3 J sin^2 J = cos^3 J - cos^5 J

Answer by lwsshak3(11628)

Trying to prove this is an identity.
cos^3 J sin^2 J = cos^3 J - cos^5 J
**
Starting with left side
cos^3 J sin^2 J=cos^3 J (1-cos^2 J)=cos^3 J - cos^5 J
verified:
left side=right side